# Quantum theory of surface lattice resonances

### v2

@@ Abstract
The collective interactions of nanoparticles arranged in periodic structures give rise to high-$Q$ in-plane diffractive modes known as surface lattice resonances (SLRs). While these resonances and their broader implications have been extensively studied within the framework of classical electrodynamics and linear response theory, a quantum optical theory capable of describing the dynamics of these structures, especially in the presence of material nonlinearities, beyond the use of ad hoc few-mode approximations is largely missing. To this end, we consider here a lattice of metallic nanoparticles coupled to the background electromagnetic field and derive the quantum input-output relations within the electric dipole approximation. As a first application, we discuss coupling of the nanoparticle array to an external array of emitters and illustrate how the formalism can be extended to molecular optomechanics. Specifically, we show how the high $Q$-factors provided by SLRs can facilitate coupling between the nanoparticle array and collective vibrational modes. Finally, we examine a scenario in which the metallic nanoparticles themselves are replaced with saturable excitonic emitters, illustrating how the nonlinearity of the emitters can be used to switch the SLR condition between different electronic transitions. We further show, using a perturbative approach that explicitly accounts for population dynamics, how the resulting switching dynamics can be observed and characterized in pump–probe experiments. Our work provides a seamless framework to rigorously model SLR phenomena with quantum emitters without resorting to phenomenological treatments of either the electromagnetic environment or the material degrees of freedom.
@@ VII Acknowledgments
S3. Optomechanical interaction
@@
$\dot{\hat{b}}_{j}$ $=-(\mathrm{i}\omega_{\mathrm{vib}}+\Gamma_{\mathrm{vib}})\hat{b}_{j}-\mathrm{i}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}},$ (S38c)
@@
$\dot{\hat{b}}_{j}^{\dagger}$ $=\left[\mathrm{i}(\omega_{\mathrm{vib}}-\omega_{\ell})-\Gamma_{\mathrm{vib}}\right]\hat{b}_{j}^{\dagger}+\mathrm{i}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}},$ (S45c)
@@
S4. Nonlinear polarizability of excitonic SLRs for static populations
@@
S5. Pump-probe spectroscopy of nonlinear SLR dynamics
@@
where the $\ldots$ denotes other terms which become important beyond third order for our perturbative treatment below (third order coherences feeding back into populations) and which we therefore drop for the ease of notation.
@@
$\langle\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\rangle$ $=\eta_{p}f_{p}(t)\,\mathrm{e}^{-\mathrm{i}\omega_{p}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},\qquad\langle\hat{\sigma}^{23,\mathrm{in}}_{j}(t)\rangle=\eta_{p^{\prime}}f_{p^{\prime}}(t-\tau_{\Delta})\,\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel}\cdot\mathbf{r}_{j}},$ (S54)
where $\eta_{p,p^{\prime}}$ describes the pulse amplitudes (considered real here), $f_{p,p^{\prime}}(t)$ the temporal envelope (assumed Gaussian), $\tau_{\Delta}$ is the delay time between the pulses, and we assume collinear incidence of pump and probe pulses at wavevector $\mathbf{k}_{\parallel}$ parallel to the array plane. We then go to momentum space for both coherences and population operators
@@
At each order, a closed set of equations for the expansion coefficients $O_{\mathbf{q}}^{(n)(m)}$ can then be derived. In particular, we will consider the modification of the probe response due to the second-order populations generated by the pump, corresponding to order $(2)(1)$ [see Feynman diagrams in Fig. S2].
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}-\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}p_{\mathrm{inv}}^{12}(0)-\delta_{\mathbf{q},\mathbf{k}_{\parallel}}f_{p}(\omega-\omega_{p})p_{\mathrm{inv}}^{12}(0),$ (S57)
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\delta_{\mathbf{q},\mathbf{k}_{\parallel}}f_{p}(\omega-\omega_{p}).$ (S58)
Therefore, at linear order only the momentum mode $\mathbf{q}=\mathbf{k}_{\ell,\parallel}$ is driven (and $\mathbf{q}=-\mathbf{k}_{\ell,\parallel}$ for the complex conjugate coherence $\sigma_{\mathbf{q}}^{12,*}$).
@@
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}}(\omega)^{(2)(0)}$ $=-2\mathrm{Im}\sum_{\mathbf{p}}\left[\sigma_{\mathbf{p}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{q}-\mathbf{p}}^{12}\sigma_{\mathbf{q}-\mathbf{p}}^{12}(\omega)^{(1)(0)}\right]$ (S59)
$+\left[\sigma_{\mathbf{q}-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{q}+\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right],$
generally leading to convolutions in both momentum and frequency space. Here, an asterisk in the exponent describes complex conjugation and an asterisk between two symbols denotes convolution. Making use of the fact that the only momentum mode driven for the coherence at linear order is $\mathbf{k}_{\parallel}$ (and the opposite sign for the complex conjugate), we can see that only the zero-momentum mode is driven for the population at second order (see Fig. S2)
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}=0}(\omega)^{(2)(0)}$ $=-2\mathrm{Im}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{k}_{\parallel}}^{12}\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$ (S60)
$+\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right].$
In analogy to Sec. S1, we pull the self-interaction term at $\mathbf{r}_{\Lambda}=0$ out of the lattice sum $\bar{\mathcal{S}}_{\mathbf{q}}^{12}$ [see Eq. (S23)], giving rise to radiative damping for the population:
$-2\mathrm{Im}\left[\left(\frac{\mathrm{i}\Gamma_{\mathrm{rad}}^{12}}{2}\right)\left(\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right)\right]$ $=-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\mathrm{Re}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$
$-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right],$ (S61)
where the last step holds since $\sigma^{*}\ast\sigma$ is the Fourier transform of a real time‑domain quantity and hence has Hermitian symmetry.
Therefore, in total we obtain for the second order populations
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}=0}(\omega)^{(2)(0)}$ $=-\Gamma_{\mathrm{12}}^{\mathrm{rad}}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,\ast}(\omega)^{(1)(0)}\ast\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]-2\mathrm{Im}\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast\bar{\mathcal{S}}_{\mathbf{k}_{\parallel}}^{12}\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\right]$ (S62)
$+\left[\sigma_{-\mathbf{k}_{\parallel}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{k}_{\parallel}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right].$
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}$ $=-\left(\mathrm{i}\omega_{23}+\frac{\Gamma_{23}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(2)(1)}+\left[f_{p^{\prime}}(\omega-\omega_{p^{\prime}})\ast\left(\sigma^{12,\dagger}\sigma^{12}\right)_{\mathbf{q}-\mathbf{k}_{\parallel}}(\omega)^{(2)(0)}\right](\omega)+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{23}(\omega)\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}.$ (S63)
Therefore, the momentum mode that is driven for the coherence at third order is $\mathbf{q}=\mathbf{k}_{\parallel}$, since only the population with zero quasi-momentum is contributing.

### v3

@@ Abstract
The collective interactions of nanoparticles arranged in periodic structures give rise to high-$Q$ in-plane diffractive modes known as surface lattice resonances. While these resonances and their broader implications have been extensively studied within the framework of classical electrodynamics and linear response theory, a quantum optical theory capable of describing the dynamics of these structures, especially in the presence of material nonlinearities beyond ad hoc few-mode approximations, is largely missing. To this end, we consider a lattice of metallic nanoparticles coupled to the electromagnetic field and derive the quantum input–output relations within the electric dipole approximation. As applications, we analyze coupling between the nanoparticle array and external quantum emitters, and show how the formalism extends to molecular optomechanics, where the high $Q$-factors of SLRs enable coupling to collective vibrational modes. We further consider arrays composed of saturable excitonic emitters, demonstrating how emitter nonlinearities can be used to switch the SLR condition between electronic transitions. Using a perturbative approach that accounts for population dynamics, we show how these effects can be probed in pump–probe experiments and give rise to nonlinear phase-matching phenomena. Our work provides a microscopic framework for modeling SLRs interacting with quantum emitters without phenomenological descriptions of the electromagnetic environment.
@@ VII Acknowledgments
S3. Extinciton spectrum for different lattice constants
Figure S2: Extinction spectrum for different lattice constants. Extinction spectra $\sigma_{\mathrm{ext}}(\mathbf{k})$ (top) as well as cross sections (illustrated by the green dashed curves) at normal incidence (bottom) for lattice constants (a) $a=520\,\mathrm{nm}$, (b) $a=550\,\mathrm{nm}$, and (c) $a=580\,\mathrm{nm}$ for a resonance wavelength of $\lambda_{0}=500\,\mathrm{nm}$ and radiative linewidth $\Gamma_{0}=0.5\,\mathrm{eV}$. As the lattice constant increases, the Rayleigh anomaly shifts to longer wavelength relative to the particle resonance, resulting in a progressive narrowing of the SLR linewidth.
S4. Optomechanical interaction
@@
$\dot{\hat{b}}_{j}$ $=-(\mathrm{i}\omega_{\mathrm{vib}}+\Gamma_{\mathrm{vib}}/2)\hat{b}_{j}-\mathrm{i}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}},$ (S38c)
@@
$\dot{\hat{b}}_{j}^{\dagger}$ $=\left[\mathrm{i}(\omega_{\mathrm{vib}}-\omega_{\ell})-\Gamma_{\mathrm{vib}}{\color[rgb]{0,0,0}/2}\right]\hat{b}_{j}^{\dagger}+\mathrm{i}\sum_{\mathbf{k},\lambda}g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}},$ (S45c)
@@
We also note that fluctuations of molecular positions and phases within the unit cells can be incorporated by replacing $\mathbf{r}_{\mathrm{m}}\to\mathbf{r}_{\mathrm{m}}+\delta\mathbf{r}_{\mathrm{m},\Lambda}$, and $\mathbf{p}\to\mathbf{p}\mathrm{e}^{\mathrm{i}\Phi_{\Lambda}}$, which modifies the sum to
$\mathcal{S}_{\mathbf{q}}^{\mathrm{OM}}(\omega)=\frac{\mu_{0}|\mathbf{p}|\omega^{2}}{c^{2}\hbar\epsilon_{0}}\sum_{\Lambda}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot(\mathbf{r}_{\Lambda}+\mathbf{r}_{\mathrm{m}}+\delta\mathbf{r}_{\mathrm{m},\Lambda})}\mathrm{e}^{\mathrm{i}\Phi_{\Lambda}}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\cdot\mathbf{G}(\mathbf{r}_{\Lambda}+\mathbf{r}_{\mathrm{m}}+\delta\mathbf{r}_{\mathrm{m},\Lambda},\omega)\cdot\boldsymbol{\varepsilon}_{\mathbf{p}},$ (S47)
generally leading to a renormalization of the lattice sum and therefore the optomechanical coupling. Since for most of the results presented in the main text, we focus on modes with $|\mathbf{q}|\approx 0$, whose wavelength is much larger than the lattice constant. In this long-wavelength limit, the factors $\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot\delta\mathbf{r}_{\mathrm{m},\Lambda}}$ are close to unity, such that fluctuations of the molecular positions (and similarly of the phases) are strongly suppressed in the lattice sum. As a result, these non-idealities produce only minor corrections and do not modify the qualitative behaviour discussed in the main text.
S5. Nonlinear polarizability of excitonic SLRs for static populations
@@
S6. Pump-probe spectroscopy of nonlinear SLR dynamics
@@
where the $\ldots$ denotes other terms which become important beyond third order for our perturbative treatment below (third-order coherences feeding back into populations) and which we therefore drop for the ease of notation.
@@
$\langle\hat{\sigma}_{j}^{12,\mathrm{in}}(t)\rangle$ $=\eta_{p}f_{p}(t)\,\mathrm{e}^{-\mathrm{i}\omega_{p}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel,p}\cdot\mathbf{r}_{j}},\qquad\langle\hat{\sigma}^{23,\mathrm{in}}_{j}(t)\rangle=\eta_{p^{\prime}}f_{p^{\prime}}(t-\tau_{\Delta})\,\mathrm{e}^{-\mathrm{i}\omega_{p^{\prime}}t}\mathrm{e}^{\mathrm{i}\mathbf{k}_{\parallel,p^{\prime}}\cdot\mathbf{r}_{j}},$ (S55)
where $\eta_{p,p^{\prime}}$ describes the pulse amplitudes (considered real here), $f_{p,p^{\prime}}(t)$ the temporal envelope (assumed Gaussian), $\tau_{\Delta}$ is the delay time between the pulses, and we assume incidence of pump and probe pulses at wavevectors $\mathbf{k}_{\parallel,p/p^{\prime}}$. We then go to momentum space for both coherences and population operators
@@
At each order, a closed set of equations for the expansion coefficients $O_{\mathbf{q}}^{(n)(m)}$ can then be derived under the mean-field assumption $\expectationvalue{\hat{\sigma}_{j}\hat{\sigma}_{j^{\prime}}}=\expectationvalue{\hat{\sigma}_{j}}\expectationvalue{\hat{\sigma}_{j^{\prime}}}=\sigma_{j}\sigma_{j^{\prime}}$ for $j\neq j^{\prime}$ (and similarly for $\expectationvalue{\hat{p}_{\mathrm{inv},j}\hat{\sigma}_{j^{\prime}}}$). In particular, in the following we will consider the modification of the probe response due to the second-order populations generated by the pump, corresponding to order $(2)(1)$ [see Feynman diagrams in Fig. S3].
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}-\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}p_{\mathrm{inv}}^{12}(0)-\delta_{\mathbf{q},\mathbf{k}_{\parallel,p}}f_{p}(\omega-\omega_{p})p_{\mathrm{inv}}^{12}(0),$ (S58)
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}$ $=-\left(\mathrm{i}\omega_{12}+\frac{\Gamma_{12}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{12}(\omega)\sigma_{\mathbf{q}}^{12}(\omega)^{(1)(0)}+\delta_{\mathbf{q},\mathbf{k}_{\parallel,p}}f_{p}(\omega-\omega_{p}).$ (S59)
Therefore, at linear order only the momentum mode $\mathbf{q}=\mathbf{k}_{\parallel,p}$ (we restrict all wavevectors to the 11st Brillouin zone) is driven (and $\mathbf{q}=-\mathbf{k}_{\parallel,p}$ for the complex conjugate coherence $\sigma_{\mathbf{q}}^{12,*}$).
@@
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}}(\omega)^{(2)(0)}$ $=-\Gamma_{12}^{\mathrm{rad}}\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}}(\omega)^{(2)(0)}-2\mathrm{Im}\sum_{\mathbf{p}}\left[\sigma_{\mathbf{p}}^{12,*}(\omega)^{(1)(0)}\ast\mathcal{S}_{\mathbf{q}-\mathbf{p}}^{12}\sigma_{\mathbf{q}-\mathbf{p}}^{12}(\omega)^{(1)(0)}\right]$ (S60)
$+\left[\sigma_{\mathbf{q}-\mathbf{k}_{\parallel,p}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{q}+\mathbf{k}_{\parallel,p}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right],$
generally leading to convolutions in both momentum and frequency space. Here, an asterisk in the exponent describes complex conjugation and an asterisk between two symbols denotes convolution. Making use of the fact that the only momentum mode driven for the coherence at linear order is $\mathbf{k}_{\parallel,p}$ (and the opposite sign for the complex conjugate), we can see that only the zero-momentum mode is driven for the population at second order (see Fig. S3)
$-\mathrm{i}\omega\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}=0}(\omega)^{(2)(0)}$ $=-\Gamma_{12}^{\mathrm{rad}}\left[\sigma^{12,\dagger}\sigma^{12}\right]_{\mathbf{q}}(\omega)^{(2)(0)}-2\mathrm{Im}\left[\sigma_{-\mathbf{k}_{\parallel,p}}^{12,*}(\omega)^{(1)(0)}\ast\mathcal{S}_{\mathbf{k}_{\parallel,p}}^{12}\sigma_{\mathbf{k}_{\parallel,p}}^{12}(\omega)^{(1)(0)}\right]$ (S61)
$+\left[\sigma_{-\mathbf{k}_{\parallel,p}}^{12,*}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right]+\left[\sigma_{\mathbf{k}_{\parallel,p}}^{12}(\omega)^{(1)(0)}\ast f(\omega-\omega_{p})\right].$
@@
$-\mathrm{i}\omega\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}$ $=-\left(\mathrm{i}\omega_{23}+\frac{\Gamma_{23}^{\mathrm{rad}}}{2}\right)\sigma_{\mathbf{q}}^{12}(\omega)^{(2)(1)}+\left[f_{p^{\prime}}(\omega-\omega_{p^{\prime}})\ast\left(\sigma^{12,\dagger}\sigma^{12}\right)_{\mathbf{q}-\mathbf{k}_{\parallel,p^{\prime}}}(\omega)^{(2)(0)}\right](\omega)+\mathrm{i}\mathcal{S}_{\mathbf{q}}^{23}(\omega)\sigma_{\mathbf{q}}^{23}(\omega)^{(2)(1)}.$ (S62)
Therefore, the momentum mode that is driven for the coherence at third order is $\mathbf{q}=\mathbf{k}_{\parallel,p^{\prime}}$, since only the population with zero quasi-momentum is contributing.